# Beyond Trial-and-Error: Predicting User Abandonment After a Moderation Intervention

### v2

@@ Abstract
Current content moderation practices follow the trial-and-error approach, meaning that moderators apply sequences of interventions until they obtain the desired outcome. However, being able to preemptively estimate the effects of an intervention would allow moderators the unprecedented opportunity to plan their actions ahead of application. As a first step towards this goal, here we propose and tackle the novel task of predicting the effect of a moderation intervention. We study the reactions of 16,540 users to a massive ban of online communities on Reddit, training a set of binary classifiers to identify those users who would abandon the platform after the intervention—a problem of great practical relevance. We leverage a dataset of 13.8M posts to compute a large and diverse set of 142 features, which convey information about the activity, toxicity, relations, and writing style of the users. We obtain promising results, with the best-performing model achieving micro F1 $=0.800$ and macro F1 $=0.676$. Our model demonstrates robust generalizability when applied to users from previously unseen communities. Furthermore, we identify activity features as the most informative predictors, followed by relational and toxicity features, while writing style features exhibit limited utility. Our results demonstrate the feasibility of predicting the effects of a moderation intervention, paving the way for a new research direction in predictive content moderation aimed at empowering moderators with intelligent tools to plan ahead their actions.
Index Terms:
Content moderation, predictive moderation, user abandonment, machine learning.
I Introduction
Online social media has brought about the sharing of opinions and information of all sorts, locally and across the globe [1]. However, this ease of communication has also caused multiple problems, including the spread of toxic content, which is not only capable of causing harm to those users targeted by such misbehavior, but also to the overall health and inclusivity of the online environments [2]. Online platforms are thus required to enforce content moderation actions to mitigate this and other issues [3]. Moderation strategies can be broadly divided into two main categories: hard and soft. The former consists of permanently removing users, groups, or posts that are violating community guidelines, such as posts deemed toxic. This type of intervention —also called deplatforming— can have unintended consequences because certain users perceive content removals as a form of censorship, especially if unmotivated. As a consequence, they may feel less prone to posting again, or they could even decide to migrate to other platforms in which aggressive behavior is tolerated [4, 5, 6]. This led to the proposal of the second broad form of moderation —soft moderation— which warns users about the harmfulness of a content without removing it [7, 8]. An example of a soft moderation intervention is Reddit’s quarantine, which affects the visibility of a whole subreddit (i.e., a community within Reddit) by limiting access to it and by warning new users about the potential harmful content therein [9].
@@ I Introduction
The growing consensus among the thriving content moderation literature is that gradual approaches that initially apply less punitive interventions, and that subsequently escalate the severity whenever required, are generally more effective than other solutions [10]. Nevertheless, hard moderation is still the most widely used form of content moderation. A well-known example is the so-called Great Ban, a massive deplatforming intervention enforced by Reddit in response to the rise of toxic and hateful speech on the platform, which involved the permanent ban of around 2,000 subreddits in June 2020 [11]. Among the banned subreddits was r/The_Donald, a community of Trump supporters that was one of the most popular political subreddits at the time [12]. Before the ban the community had already faced two milder moderation interventions: a quarantine and a restriction that limited its visibility and the users that could serve as moderators of the community [13]. This sequence of interventions has been the subject of several studies that assessed changes in activity levels, use of language, political bias, factual news sharing, and toxicity after the interventions [13, 14, 15, 16, 17]. Interestingly, these studies revealed varying user reactions to the interventions. While some users exhibited a decrease in toxicity, others increased it, and in some instances very noticeably [14]. Many users were not much affected by the interventions, showing no significant change in behavior afterward. Others, however, seemed to be resentful of the platform and significantly decreased their activity, so much so that some completely abandoned the platform [11].
@@
The previous results highlight the complexity of content moderation. First, in content moderation one size does not fit all, and effective moderation interventions should necessarily be tailored to the targets of the moderation [18]. Furthermore, the current moderation strategy follows a trial-and-error approach, where platforms apply sequences of interventions until they meet the desired outcome, or force the misbehaving users out of the platform. r/The_Donald was a prime example of this strategy, as it was subject to two unsuccessful interventions prior to the Great Ban. Given that the first of such interventions was applied in June 2019, while the last in June 2020, it is evident that it took several months for the platform to evaluate the effects of each intervention and to impose further restrictions. Importantly, during that time the misbehaving users continued to spread toxic content and to engage in harmful behaviors.
@@
There is thus a pressing need for social media platforms to more efficiently evaluate the impact of a potential moderation intervention before its application. This would open up the possibility to plan interventions ahead of time, rather than to assess and correct afterwards. To date, however, no evidence exists as to whether the effects of a moderation intervention are predictable.
I-A Contributions
@@
Motivated by the need for a strategic planning and deployment of moderation actions, we tackle the new task of predicting the effect of a moderation intervention. While many studies carried out post-hoc descriptive analyses of the effects of moderation interventions [15, 8, 16], here for the first time we adopt a predictive approach. We leverage the results of previous studies on the Great Ban as our ground-truth for the effects of that intervention [19, 11]. Then, we analyze user behavioral traces before the Great Ban to predict possible changes in the behavior of those users after the ban. We specifically focus on predicting those users who abandon the platform after the ban, in a binary classification task. This choice is both theoretically and practically relevant. Changes in user activity are among the most widely studied effects of past moderation interventions [13, 19, 9]. As such, from the theoretical standpoint it is interesting to assess the extent to which those effects are predictable. Additionally, the task is also practically relevant since the loss of users negatively affects platform revenues, as user engagement and popularity are paramount in the economic success of social media [20, 21]. Moreover, the abandonment of influential users can also trigger chain reactions, as their followers may also choose to leave the platform, further worsening the issue. The main results of this work are summarized in the following:
@@
  • We evaluated different classification algorithms, with different sets of features and hyper-parameters, for the new task of predicting the effects of the Great Ban. Our best classifier achieved a promising micro F1 $=0.800$.
@@
I-B Significance
This work takes the first step towards providing strategic evidence and tools to support online moderators in planning their interventions. Our promising —yet preliminary— results validate the adoption of predictive models for assessing the outcomes of moderation interventions. This new approach marks a significant advancement in content moderation, offering moderators and platform administrators the ability to strategically deploy interventions to minimize adverse effects such as user attrition to less regulated platforms, while concurrently maximizing desired outcomes such as the reduction of toxic content. By enabling proactive decision-making based on predicted intervention effects, this approach stands to enhance the efficacy and efficiency of content moderation in fostering healthier online environments.
II Related Works
While many works in the literature have already studied the effects of moderation interventions from a descriptive point of view, fewer have covered predictive modeling. In this section, we summarize and discuss previous works from both perspectives.
II-A Descriptive moderation
@@
In discussing the many works that took a descriptive approach to analyzing moderation interventions, we first survey studies on the Great Ban and other hard interventions, because they are the most similar to our present work. Among the works that studied the Great Ban is [19], which described the changes in activity and in-group vocabulary of those users who participated in the 15 most popular subreddits out of the 2,000 shut with the ban. The results highlighted the heterogeneity of the effects of the intervention and that top users tended to reduce their activity and their use of in-group vocabulary. The work in [11] carried out a similar analysis but focused on evaluating changes in toxicity rather than activity. Nonetheless the authors found that 15.6% of the affected users abandoned Reddit after the ban. Among those who stayed there was a general reduction in toxic comments, but a small subset of users drastically increased their toxicity. Other works that evaluated the effects of hard content moderation include [22, 16]. The former measured changes occurred in Google Trends and on Wikipedia, while the latter on Twitter. Both works revealed that banning toxic influencers reduced the attention toward them and the number of posts citing them. These studies concluded that deplatforming may be a valid moderation strategy, when used appropriately.
@@
Due to its prominence on Reddit and the multiple interventions that it faced, r/The_Donald was the focus of many works that investigated the effectiveness of online content moderation. For example, some works showed how the quarantine was able to reduce the visibility of the subreddit. At the same time however, it was unsuccessful at reducing racist and misogynistic language [9, 15]. Others examined the interventions in terms of activity, toxicity, factual reporting, and political bias both at community and user levels [13, 14]. Community-level results found a general reduction in activity, a strong long-term increase in toxicity, a slight decrease in factual reporting, and no particular change in political ideology [13]. However, user-level reactions were more diversified and sometimes even extreme, especially for the most active users [14]. These results highlight that community-level effects are not always representative of the underlying user-level effects, which once again reaffirms the limitations of one-size-fits-all moderation [18]. On the contrary, studying, and possibly even predicting user-level reactions to moderation interventions could be particularly beneficial to moderators.
II-B Predictive moderation
@@
Recent computational efforts in the study of moderation interventions increasingly adopt predictive rather than descriptive approaches. Among this body of work, the vast majority of studies tackled the task of predicting which pieces of content would be subject to moderation. For example, the authors of [23] predicted with high accuracy which YouTube videos would be later moderated. The predictions were accurate even at posting-time, reducing the disappointment users face upon deletion after publication. Similarly, [24] proposed Lambretta, a learning to rank system developed to identify tweets likely to be removed from the platform because they convey false information. The system is intended to provide support to moderators in preventing the spread of fake news. Along the same line, CrossMod [25] automatically detects which Reddit posts are likely to be removed by the moderators. The system is also capable of taking actions based on certain conditions set by the moderators themselves. The aforementioned works serve a twofold goal. On the one hand, they contribute to reverse engineer the content moderation actions taken by large online platforms, which often lack transparency and consistency [3]. On the other hand, the automated systems developed as part of the studies could also support human moderators by reducing the burden of extensive manual analyses as well as by reducing the emotional toll that comes from being continuously exposed to toxic, harmful, or otherwise inappropriate content [26].
Other related works focused on predicting user behaviors, such as [27] that aimed to identify those users that would evade a Wikipedia ban by creating a brand new account. Additional tasks in the same work included early detection of such evasion and account matching. Instead, [28] discussed the feasibility of proactive moderation interventions on Reddit. Results showed that Reddit communities are constantly evolving, and that communities bound to become toxic can be preemptively identified. This early detection opens up the possibility to intervene before the problem escalates, thus possibly resulting in the application of less restrictive and punitive interventions.
@@
This survey of the existing literature in predictive content moderation showed that all previous works focused on detecting which content, users, or communities would later face or evade moderation. Conversely, to the best of our knowledge, no one has ever focused on predicting the outcome of a moderation intervention. Our present work contributes to filling this gap.
III Problem Definition
@@
We introduce the new task of predicting the dichotomous effect on user abandonment (yes or no) of a moderation intervention on a social media community. We first define the general framing of the problem, then we specify our binary classification task in the context of Reddit’s Great Ban.
@@
III-A Problem framing
Let $i$ be a moderation intervention that occurred at time $t_{0}$, for which we delimit a surrounding time frame of interest that starts at $t_{-m}$ and ends at $t_{+n}$. This time frame is thus composed of a pre-intervention period $P^{-}$ that covers $[t_{-m},t_{0})$, and a post-intervention period $P^{+}$ that covers $(t_{0},t_{+n}]$. We exclude $t_{0}$ at which $i$ occurred because it is likely that user activity at this time is non-representative due to the application of the intervention itself (e.g., immediate and reactionary spikes in activity). Let $P^{*}$ be the abandonment time window after the intervention that starts at offset $t_{+a}$ and ends at $t_{+n}$, such that $P^{*}\subseteq P^{+}$. If there is no activity for a given user within the community of interest during $P^{*}$, we consider that user as abandoning. We also consider that the closer the abandonment offset $t_{+a}$ to $t_{0}$, the harder is the abandonment, and the farther the softer, as illustrated in Fig. 1. For instance, in the case of $P^{*}=P^{+}$ with a user having non-zero activity immediately after the intervention and then no activity for the rest of $P^{*}$, the user would be considered as non-abandoning due to the rigidness of the $t_{+a}=t_{0}$ offset. However, it could be argued that users with no activity levels for several consecutive months some time after the intervention can hardly be considered active even if they had some activity immediately after the intervention. Hence, the selection of a larger offset $t_{+a}\gg t_{0}$ softens the abandonment time window $P^{*}$. At the same time, $t_{+a}$ must not be too close to $t_{+n}$, otherwise $P^{*}$ would be too short to be useful.
The classification task $C$ is thus defined as predicting those users who were active in a given community before the intervention during $P^{-}$, but not afterward during the time window $P^{*}$, based on a dataset of user features $D^{-}$ from $P^{-}$. This dataset is defined as:
$D^{-}=\{(X_{1},Y_{1},\ldots,Z_{1}),\ldots,(X_{N},Y_{N},\ldots,Z_{N})\}$
where $X_{u},Y_{u},\ldots,Z_{u}$ are the values of the variables chosen as predictive features during $P^{-}$ for a given user $u$.
In other words, let $A^{-}$ be the activity level of user $u$ during $P^{-}$ and $A^{*}$ the activity level for $u$ during $P^{*}$. We then assign each user $u$ to one of the following two classes:
$C_{-1}$ =$=\{u|A^{-}>0\;\text{and}\;A^{*}>0\}$
$C_{+1}$ =$=\{u|A^{-}>0\;\text{and}\;A^{*}=0\}$
with the negative class $C_{-1}$ identifying non-abandoning users and the positive class $C_{+1}$ identifying abandoning users. Then, the objective of the abandonment task is to learn the function $f$ that, based on dataset $D^{-}$, assigns the correct class to each user $u$:
$f(X_{u},Y_{u},\ldots,Z_{u})\approx C_{u}\in\{C_{-1},C_{+1}\}.$ (1)
III-B Task specification

### v3

# 1]organization=IIT-CNR, addressline=Via Moruzzi 1, code=56124, city=Pisa, country=Italy

@@ Abstract
Current content moderation follows a reactive, trial-and-error approach, where interventions are applied and their effects are only measured post-hoc. In contrast, we introduce a proactive, predictive approach that enables moderators to anticipate the impact of their actions before implementation. We propose and tackle the new task of predicting user abandonment following a moderation intervention. We study the reactions of 16,540 users to a massive ban of online communities on Reddit, training a set of binary classifiers to identify those users who would abandon the platform after the intervention—a problem of great practical relevance. We leverage a dataset of 13.8 million posts to compute a large and diverse set of 142 features, which convey information about the activity, toxicity, relations, and writing style of the users. We obtain promising results, with the best-performing model achieving micro F1-score $=0.914$. Our model shows robust generalizability when applied to users from previously unseen communities. Furthermore, we identify activity features as the most informative predictors, followed by relational and toxicity features, while writing style features exhibit limited utility. Theoretically, our results demonstrate the feasibility of adopting a predictive machine learning approach to estimate the effects of moderation interventions. Practically, this work marks a fundamental shift from reactive to predictive moderation, equipping platform administrators with intelligent tools to strategically plan interventions, minimize unintended consequences, and optimize user engagement.
keywords:
[…]
1 Introduction
@@
Online social media has greatly facilitated the sharing of information and opinions of all sorts, both locally and across the globe [1]. However, this also includes less desirable behavior, such as the spread of online toxic content, which besides potentially harming targeted users, is detrimental to the overall health and inclusivity of online environments [2, 3]. Online platforms are thus constrained to moderate such content, not only to maintain a safe and welcoming environment, but also because of legal compliance requirements [4]. Moderation strategies can be broadly divided into two main categories: hard and soft. The former consists of permanently removing users, groups, or posts that are violating community guidelines, such as posts deemed toxic. This type of intervention —also called deplatforming— can have unintended consequences because certain users perceive content removals as a form of censorship, especially if unmotivated. As a consequence, they may feel less prone to post again, or they could even decide to migrate to other more tolerant platforms tolerated [5, 6, 7]. This led to the proposal of the second broad form of moderation —soft moderation— which warns users about the harmfulness of a content without removing it [8, 9]. The growing consensus among the thriving content moderation literature is that gradual approaches that initially apply less punitive interventions, and that subsequently escalate the severity whenever required, are generally more effective than other solutions [10]. Nevertheless, hard interventions are still the most widely used form of content moderation. A well-known example of this kind is the so-called Great Ban, a massive deplatforming intervention enforced by Reddit in response to the rise of toxic and hateful speech on the platform, which involved the permanent ban of around 2,000 subreddits in June 2020 [11]. Among the banned subreddits was r/The_Donald, a community of Trump supporters that was one of the most popular political subreddits at the time [12]. Before the ban, the community had already faced two milder moderation interventions: a quarantine and a restriction that limited its visibility and the users that could serve as moderators of the community [13]. This sequence of interventions has been the subject of several studies that assessed changes in activity levels, use of language, political bias, factual news sharing, and toxicity [13, 14, 15, 16, 17]. These works found that many users were not much affected by the interventions, showing no significant change in behavior afterward. However, others seemed to be resentful of the platform and significantly decreased their activity, so much so that some completely abandoned the platform [11, 18].
@@
The previous results highlight the complexity of content moderation. First, in content moderation one size does not fit all, and effective moderation interventions should necessarily be tailored to the targets of the moderation [19, 20]. Furthermore, the current moderation strategy follows a trial-and-error approach, where platforms apply sequences of interventions until they meet the desired outcome, or force the misbehaving users out of the platform. r/The_Donald was a prime example of this strategy, as it was subject to two unsuccessful interventions prior to the Great Ban [13]. Given that the first of such interventions was applied in June 2019, while the last in June 2020, it took several months for the platform to evaluate the effects of each intervention and to impose further restrictions. Importantly, during that time the misbehaving users continued to spread toxic content and to engage in harmful behaviors.
@@
There is thus a pressing need to develop intelligent systems capable of supporting moderators and platform administrators by estimating the likely effect of a moderation intervention before its application. This would open up the possibility to plan interventions ahead of time, rather than to assess and correct afterwards. To date, however, no evidence exists as to whether the effects of a moderation intervention are predictable.
1.1 Contributions
@@
Motivated by the need for a strategic planning and deployment of moderation actions, we introduce and tackle the new task of predicting the effect of a moderation intervention. While many studies carried out post-hoc descriptive analyses of the effects of moderation interventions [15, 9, 16], here for the first time we adopt a predictive approach. We leverage the results of previous studies on the Great Ban as our ground-truth for the effects of that intervention [21, 11]. Then, we analyze user behavioral traces before the Great Ban to predict possible changes in the behavior of those users after the ban. We specifically focus on predicting those users who abandon the platform after the ban, in a binary classification task. This choice is both theoretically and practically relevant. Changes in user activity are among the most widely studied effects of past moderation interventions [13, 21, 22]. As such, from the theoretical standpoint it is interesting to assess the extent to which those effects are predictable. Additionally, the task is also practically relevant since the loss of users negatively affects platform revenues, as user engagement and popularity are paramount in the economic success of social media [23, 24]. Moreover, the abandonment of influential users can also trigger chain reactions, as their followers may also choose to leave the platform, further worsening the issue. The main results of this work are summarized in the following:
@@
  • We proposed the new task of predicting users who abandon the platform following a moderation intervention and we evaluated different classification algorithms, hyper-parameters, and features sets for the task. Our best classifier is capable of detecting abandoning and non-abandoning users with a promising micro F1 score $=0.914$.
@@
1.2 Significance and Applicability
Our predictive approach enhances the moderation process by enabling platform administrators to anticipate the effects of moderation interventions, allowing more informed and effective decision-making on a tighter feedback loop, as depicted in Figure 1. In our specific case, administrators can estimate the likelihood of user abandonment after a large-scale intervention and tailor their moderation strategy accordingly. Interventions such as the collective ban of subreddits, analyzed in this study, are relatively infrequent but not uncommon. These target entire groups and communities rather than individual users or posts, usually at the behest of high-level platform administrators. Consequently, in this case the task does not have stringent real-time constraints, making it well-suited for periodic, data-driven assessments of potential outcomes of different interventions in terms of user abandonment. Moreover, the proposed approach is broadly applicable to other platforms as well, provided that historical data on past interventions is available. For instance, on Facebook, it could help predict the effects of shutting down certain groups or pages, while on Twitter/X, a similar approach could estimate the impact of banning high-profile influencers with large follower bases. Further, on the long term, such an approach could also pave the way to predictive models apt to operate in more dynamic, real-time contexts, enabling finer-grained interventions targeting individual users or posts. This shift from reactive to proactive moderation stands to enhance both the efficacy and efficiency of platform governance, fostering healthier and more sustainable digital ecosystems.
Figure 1: Overview of the moderation process with a predictive approach. Our work on predicting user abandonment following a large-scale moderation intervention is an example of how simulating different interventions and outcomes can improve platform moderation processes by opening up the possibility to take better-informed decisions.
2 Related Works
The majority of existing literature on content moderation is based on descriptive analyses, while only a small subset of works have adopted a predictive approach—none of which have specifically addressed the task of estimating the effects of moderation interventions. This section summarizes and discusses both descriptive and predictive literature in content moderation, as well as some relevant works in the related area of churn prediction.
2.1 Descriptive moderation
@@
In discussing the many works that took a descriptive approach to analyzing moderation interventions, we first survey studies on the Great Ban and other hard interventions, because they are the most similar to our present work. Among the works that studied the Great Ban is [21], which described the changes in activity and in-group vocabulary of those users who participated in the 15 most popular subreddits out of the 2,000 shut with the ban, using a word frequency divergence. The results highlighted the heterogeneity of the effects of the intervention and that top users tended to reduce their activity and their use of in-group vocabulary. The studies in [11, 18] carried out similar analyses but focused on evaluating changes in toxicity rather than activity. The authors found that 15.6% of the affected users abandoned Reddit after the ban. Among those who stayed there was a general reduction in toxic comments, but a small subset of users drastically increased their toxicity. Other works that evaluated the effects of hard content moderation include [25, 16]. The former measured changes occurred in Google Trends and Wikipedia, using the Google Knowledge Graph identifiers to apply time-series analysis on the digital attention traces. The latter studied the deplatforming of Twitter influencers and its effects on the activity of their supporters. Both works revealed that banning toxic influencers reduced the attention toward them and the number of posts citing them. These studies concluded that deplatforming may be a valid moderation strategy when used appropriately.
@@
Due to its prominence on Reddit and the multiple interventions that it faced, r/The_Donald was the focus of many works that investigated the effectiveness of online content moderation. For example, some works showed how the quarantine was able to reduce the visibility of the subreddit. At the same time however, it was unsuccessful at reducing racist and misogynistic language [22, 15]. Others examined the interventions in terms of activity, toxicity, factual reporting, and political bias both at community and user levels, using interrupted time series (ITS) regression analysis and Bayesian structural time series (BSTS) modeling [13, 14]. Community-level results found a general reduction in activity, a strong long-term increase in toxicity, a slight decrease in factual reporting, and no particular change in political ideology [13]. However, user-level reactions were more diversified and sometimes even extreme, especially for the most active users [14]. These results highlight that community-level effects are not always representative of the underlying user-level effects, which once again reaffirms the limitations of one-size-fits-all moderation [19, 20]. On the contrary, studying, and possibly even predicting user-level reactions to moderation interventions could be particularly beneficial to moderators.
2.2 Predictive moderation
@@
Recent computational efforts in the study of content moderation increasingly adopt predictive rather than descriptive approaches. Among this body of work, the vast majority of studies tackled the task of predicting which pieces of content would be subject to moderation. For example, [26] predicted with high accuracy which YouTube videos would be later moderated using video, comment, and text-related features. The predictions were accurate even at posting-time, reducing the disappointment users face upon deletion after publication. Similarly, [27] proposed Lambretta, a learning to rank system developed to identify tweets likely to be removed from the platform because they convey false information. The system is intended to provide support to moderators in preventing the spread of fake news. Along the same line, CrossMod [28] automatically detects which Reddit posts are likely to be removed by the moderators, using cross-community learning via an ensemble of classifiers. The system is also capable of taking actions based on certain conditions set by the moderators themselves. The aforementioned works serve a twofold goal. On the one hand, they contribute to reverse engineer the content moderation actions taken by large online platforms, which often lack transparency and consistency [4]. On the other hand, the automated systems developed as part of the studies could also support human moderators by reducing the burden of extensive manual analyses as well as by reducing the emotional toll that comes from being exposed to toxic, harmful, or otherwise inappropriate content [29]. Other related works focused on predicting user behaviors, such as [30] that aimed to identify those users that would evade a Wikipedia ban by creating a new account. Additional tasks in the same work included early detection of such evasion and account matching. Instead, [31] discussed the feasibility of proactive moderation interventions on Reddit, looking at vocabulary features to identify problematic communities. Results showed that Reddit communities are constantly evolving, and that communities bound to become toxic can be preemptively identified. This early detection opens up the possibility to intervene before the problem escalates, thus possibly resulting in the application of less restrictive and punitive interventions.
@@
This survey of the existing literature on predictive content moderation showed that all previous works focused on detecting which content, users, or communities would later face or evade moderation. Conversely, to the best of our knowledge, no one has focused on predicting the outcome of a moderation intervention. Our present work contributes to filling this gap.
2.3 Churn prediction
Churn prediction is a well-studied task in customer management that aims at forecasting whether a user will stop engaging with a platform or service based on their past behavior. It is related to our novel task of predicting user abandonment following a moderation intervention since the latter can be seen as a specific case of churn influenced by platform enforcement actions. Churn prediction has been traditionally tackled through various approaches, including the use of different feature sets, data sampling techniques, and machine learning models [32]. The majority of works tackled churn prediction in the context of telecom services. Some of them leverage Long Short-Term Memory (LSTM) and Convolutional Neural Networks (CNN) [33, 34]. The former does not consider information sequence by using a Bidirectional LSTM, while the latter uses temporal features with different granularity to maximize the accuracy. Other studies in this domain addressed data skewness through balancing techniques [35] or explored model explainability [36]. Other works study user adoption and churn related to the use of web browsers [37], by using a new model called Multivariate Behavior Sequence Transformer (MBST) to explore temporal and behavioral information separately. Another set of works covers churn prediction for the usage of financial services [38, 39], the former used CNNs to analyze text, while the second work used unstructured data like customer interactions as features. Although no prior research specifically addressed user retention in social media, numerous studies have examined users lifecycles within online communities. This body of work investigated various aspects of user engagement, such as how individuals interact with platforms [40] or how the way their language changes over time could be an indicator of the platform evolution [41]. Additionally, research has explored the motivations and patterns behind user migration to alternative platforms by analyzing user comments [42]. Temporal user dynamics on social media have been a key focus in studies such as [43, 44]. The former traced user trajectories from their initial joining phase as newcomers, to their eventual departure from a platform. Instead, the latter investigated the evolution of coordinated communities over time during a campaign by evaluating user retention in single communities. Lastly, Cunha et al. highlight how variations in both user and community behaviors contribute to different levels of community success [45]
3 Problem Definition
@@
We introduce the new task of predicting whether a moderation intervention will lead a user to abandon the platform. We first outline the general framing of the problem before specifying our classification task in the context of Reddit’s Great Ban.
@@
3.1 Problem framing
Let $i$ denote a moderation intervention that takes place at time $t_{0}$. To analyze user behavior around this event, we define three key time periods, as shown in Figure 2:
  • The pre-intervention period $P^{-}=[t_{-m},t_{0})$ represents user activity before the intervention, where $t_{-m}<t_{0}$.
  • The post-intervention period $P^{+}=(t_{0},t_{+n}]$ captures user activity after the intervention, where $t_{0}<t_{+n}$.
  • The abandonment period $P^{a}=[t_{+a},t_{+n}]$, with $t_{0}<t_{+a}<t_{+n}$, is a subset of $P^{+}$ used to determine whether a user has disengaged from the platform.
Note that neither $P^{-}$ nor $P^{+}$ include $t_{0}$, since user activity at the exact moment of the intervention may be non-representative due to sudden fluctuations or spurious activity spikes. For any given user $u$, we define their activity during the aforementioned periods as follows:
  • $A_{u}^{-}$ is the pre-intervention activity of user $u$ during $P^{-}$.
  • $A_{u}^{+}$ is the post-intervention activity of user $u$ during $P^{+}$.
  • $A_{u}^{a}$ is the activity of user $u$ during the abandonment period $P^{a}$.
Using these definitions, we frame the task of predicting user abandonment as a binary classification problem, with the following two classes:
$C_{+1}$ =$=\{u\;|\;A_{u}^{-}>0\;\text{and}\;A_{u}^{a}=0\}\quad\text{({abandoning} users)}$)
[…]
The positive class $C_{+1}$ represents users who were active before the intervention but exhibited no activity during $P^{a}$, indicating abandonment. The negative class $C_{-1}$ consists of users who remained active during both $P^{-}$ and $P^{a}$, indicating continued engagement with the platform. The goal of our predictive task is to learn a function $f$ that assigns the correct class $C_{u}$ to each user $u$ based on their pre-intervention activity $A_{u}^{-}$:
$f(A_{u}^{-})\approx C_{u}\in\{C_{+1},C_{-1}\}.$ (1)
This formulation allows assessing whether pre-intervention behavior provides sufficient information to anticipate user abandonment following a moderation intervention.
Having defined the general framing of the task, we now introduce the distinction between different types of abandonment. In reality, user activity exists on a spectrum and there is no absolute boundary between full abandonment and continued participation. Some users may abruptly disengage after an intervention, others may do it gradually, while others still may reduce participation dramatically but nonetheless post sporadically. Whether certain users should be classified as abandoning or not revolves around defining a threshold for inactivity. The abandonment offset $t_{+a}$ addresses this by setting a temporal boundary: if a user exhibits no activity beyond $t_{+a}$, they are considered to have abandoned the platform. For example, this choice helps discount any brief, reactionary activity that may occur immediately after the intervention. Therefore, a shorter $t_{+a}$ captures short-term abandonment, where users disengage almost immediately following the intervention. Conversely, a longer $t_{+a}$ accounts for longer-term abandonment, recognizing users who may have remained somewhat active for a short period before eventually leaving the platform. Selecting an appropriate $t_{+a}$ provides flexibility in defining abandonment, making it possible to tailor the classification task to different analytical and practical needs.
3.2 Task specification